OpenStudy (aripotta):

If p and n are integers such that p > n > 0 and p^2 - n^2 = 12, which of the following can be the value of p - n? I. 1 II. 2 III. 4 a. I only b. II only c. I and II only d. II and III only e. I, II, and III

OpenStudy (anonymous):

\[(p+n)(p-n) = 12\]

OpenStudy (anonymous):

Firstly: \(12 = 6 \times 2\)

OpenStudy (anonymous):

\[(p+n)(p-n) = 6 \times 2\]

OpenStudy (anonymous):

If \(p=4\) and \(n=2\), then is it okay?

OpenStudy (aripotta):

wow you're a genius

OpenStudy (anonymous):

Wait..

OpenStudy (aripotta):

i wouldn't have thought to make it (p + n)(p - n)

OpenStudy (anonymous):

Above values make it, so \(p-n = 4-2 = 2\)

OpenStudy (aripotta):

right, so the answer is b?

OpenStudy (anonymous):

So, II 2, is right and we have eliminated only a option so far. :(

OpenStudy (anonymous):

Also : \(12 = 4 \times 3\), with this, we can't find \(p\) and \(n\) which can be integers.

OpenStudy (aripotta):

wait how do you know that? what does making 12 into 4 x 3 do?

OpenStudy (anonymous):

12 is product of (6*2), (4*3) and (12*1), right?? Any other set you can make?

OpenStudy (aripotta):

no, unless we do -6 and -2, etc

OpenStudy (aripotta):

but why do we have to factor the 12?

OpenStudy (anonymous):

First one ie (6*2) fits to our equation, as p =4 and n = 2 are integers and they are satisfying our equation, so p-n = 2 is sure..

OpenStudy (anonymous):

I am factoring 12, as LHS is also factored..

OpenStudy (aripotta):

oh

OpenStudy (anonymous):